OpenStudy (anonymous):

there are 32 students in a class 9 of those students are womens what percent are men

geerky42 (geerky42):

\[\dfrac{Men}{Total} = \dfrac{Total-Women}{Total}\]

OpenStudy (anonymous):

32students-9women=23men 32students.............100% 23men....................x% x=(23*100)/32 x=71,875%